OpenStudy (brucebaner):

What is the value of the expression? 4^3• 4^-5 A. 4^-15 B. 4^-2 C. 4^2 D. 4^8

OpenStudy (brucebaner):

@mathstudent55

OpenStudy (brucebaner):

@mathstudent55

OpenStudy (brucebaner):

@gerryliyana

OpenStudy (anonymous):

4^3• 4^-5 = 4^(3-5) =4^-2

OpenStudy (anonymous):

it is B

OpenStudy (brucebaner):

\[8?/8^-4=8^9\]

OpenStudy (brucebaner):

What is the missing value? A. −13 B. −5 C. 5 D. 13

OpenStudy (anonymous):

C. 5